Solve the following 5x4-x x+2 3 2 4 b. -11 F-32-17 8-10x = 5x2
Answer #1

a) 5x / 3 + ( 4 - x) / 2 = ( x+2)/4 -1

multiplying entire equation by 12

20x + 6 ( 4 - x) = 3 (x+2) - 12

20x + 24 - 6x = 3x + 6 - 12

20x - 6x - 3x = - 24 + 6 - 12

x = - 30/11

b) -11 < 5/9 F - 32 < 17

adding 32 on all sides

- 11 + 32 < 5/9 F < 17 + 32

21 < 5/9 F < 49

multiplying all sides by 9/5

189/ 5 < F < 441 / 5

c) 8 - 10x = 5x^2

5x^2 + 10x - 8 = 0

x = [ - 10 +- sqrt ( 100 + 160) ] / 10

x = -1 + sqrt 65 / 5 , -1 - sqrt 65 / 5